OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given =

OpenStudy (anonymous):



Parth (parthkohli):

Divide \(\sqrt{2}\), and I'm sorry, you had to divide by sqrt2 in the last one.

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

no this time divide what parth said

Parth (parthkohli):

sqrt2 cancels out.

OpenStudy (anonymous):

ok so then its 3?

Parth (parthkohli):

Leaves us to handle just 4.

OpenStudy (anonymous):

but how?

OpenStudy (anonymous):

:/

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {4\cancel{\sqrt{2}} \over \cancel{\sqrt{2}}} }\)

OpenStudy (anonymous):

mhmmmmmm? I think i got it i think!

OpenStudy (anonymous):

i think it was \(6\sqrt{2}\) right?

Parth (parthkohli):

Wait, is it 4sqrt2 or 6sqrt2?